OpenStudy (anonymous):

What is the volume of the cone? (Use 3.14 for π.) 355.87 ft3

jimthompson5910 (jim_thompson5910):

use the formula V = (pi*r^2*h)/3

OpenStudy (anonymous):

is my answer right or not

jimthompson5910 (jim_thompson5910):

I don't see the actual problem, so I can't confirm

jimthompson5910 (jim_thompson5910):

all I see is one potential answer, but no actual problem

OpenStudy (anonymous):

sorry hold on

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

your answer is too small

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (dan815):

she just guesses answer lol

jimthompson5910 (jim_thompson5910):

in this case r = 10 h = 17

jimthompson5910 (jim_thompson5910):

V = (pi*r^2*h)/3 V = (3.14*10^2*17)/3 V = ???

OpenStudy (anonymous):

i acally tried that one @dan815

OpenStudy (dan815):

oh ok, sry i assumed

OpenStudy (anonymous):

lol haha

OpenStudy (anonymous):

@msumner you get that wrong answer cause you did 10*2 and not 10^2

OpenStudy (anonymous):

1,067.6 ft3 ?

jimthompson5910 (jim_thompson5910):

type (3.14*10^2*17)/3 into google or wolfram alpha and they'll act as a calculator